OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 1 6x + 2y = −9

OpenStudy (anonymous):

3x + y = 1 6x + 2y = −9

OpenStudy (ghostfire):

3x+y=1 y=1-3x Then plug this in the second equation, to eliminate one variable. so you will get 6x+2(1-3x)=-9

OpenStudy (radar):

a very suspicious looking system.

OpenStudy (radar):

Note the left hand doubled but the right half didn't, there is an inconsistency here.

OpenStudy (ghostfire):

oh yeah, that's true. I didn't see that. good catch radar

OpenStudy (radar):

Sneaky to say the least, but there is an answer to inconsistency: "no answer" lol

OpenStudy (anonymous):

0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system